# Wavy Transformer

### v1

@@ Abstract
Transformers have achieved remarkable success across natural language processing (NLP) and computer vision (CV). However, deep transformer models often suffer from an over-smoothing issue, in which token representations converge to similar values as they pass through successive transformer blocks. In this paper, we establish an equivalence between the hidden-state dynamics induced by stacked attention layers and graph neural diffusion on a complete graph. From this perspective, over-smoothing can be interpreted as a consequence of the dissipative nature of the underlying diffusion dynamics. Motivated by this physical interpretation, we propose Wavy Transformer, which consists of a novel attention layer based on second-order wavy dynamics. We also introduce a feed-forward network and a normalization layer designed to preserve the physical state-velocity relationship under the chain rule, thereby extending the transformer architecture. We further validate our proposed techniques on various transformer models for NLP and CV tasks. The results consistently demonstrate that Wavy Transformer improves performance with minimal additional parameters and no extra hyperparameter tuning.
@@ 1 Introduction
Based on the above discussion, in this paper, we examine over-smoothing in transformers from the perspective of physical dynamical systems. In addition, we propose a new type of attention layer inspired by physical dynamical systems. Our approach originates from interpreting the dynamics of hidden states induced by stacked attention layers as graph neural diffusion chamberlain2021grand on a complete graph. From this viewpoint, over-smoothing can be seen as a consequence of the dissipative nature of the underlying diffusive dynamics. Motivated by this physical interpretation, we introduce Wavy Transformer, which consists of a novel attention layer based on second-order wavy dynamics. The energy-preserving property and the oscillatory behavior of the wave equation are expected to help mitigate the over-smoothing problems in transformers. We also introduce a physically inspired feed-forward network and a normalization layer designed to preserve the state-velocity relationship under the chain rule. By combining these components, we build Wavy Transformer to extend the conventional transformer architecture. Wavy Transformer block is easy to use and can be integrated into various transformer-like architecture. Furthermore, to demonstrate the effectiveness and generality of Wavy Transformer, we conduct extensive experiments on both NLP and CV tasks.
@@ 2.1 Attention and Over-smoothing
Despite their remarkable performance across a wide range of NLP and CV applications, deep transformer models often suffer from an over-smoothing problem: as the network deepens, token representations converge and become indistinguishable from one another shi2022revisiting ; wang2022antioversmooth ; nguyen2023mitigating . As a result, deeper transformers do not always outperform their shallower counterparts zhou2021deepvit . From a theoretical viewpoint, the attention layer can be understood as a low-pass filter wang2022antioversmooth , continuously erasing high-frequency information and thus reducing feature expressiveness in deeper layers. Consequently, this tendency toward over-smoothing of stacked the attention layers in deep transformer presents a critical challenge that motivates our study.
@@ 2.2 Fundamental Dynamics on Manifolds
$\frac{\partial x(u,t)}{\partial t}=\mathrm{div}\bigl{(}D\,\nabla x(u,t)\bigr{)}=D\,\Delta x(u,t),$ (3)
where $u\in M$ denotes a point (spatial coordinate) on the manifold, $t\in[0,T]$ denotes time, $D\geq 0$ is a diffusivity coefficient, $\nabla$ and $\mathrm{div}$ are respectively the gradient and divergence operators induced by the Riemannian metric on $M$, and $\Delta=\mathrm{div}\circ\nabla$ is the Laplace–Beltrami operator. Here, $D$ is assumed to be constant for simplicity. In physics, Eq.(3) models processes such as heat conduction or particle diffusion. In contrast, the hyperbolic (wave) equation on $M$ takes the form
$\frac{\partial^{2}x(u,t)}{\partial t^{2}}=\mathrm{div}\bigl{(}v^{2}\,\nabla x(u,t)\bigr{)}=v^{2}\,\Delta x(u,t),$ (4)
where $v\geq 0$ is a constant wave-propagation speed. Equation (4) governs wave-like phenomena in continuous media. We highlight the key differences between the dynamics of these PDEs.
@@ 2.2.1 Kernel Perspective of Diffusion and Wave Equations
Fundamental solutions, which are often called "kernels", represent the response to an initial impulse (e.g., a Dirac delta function). These kernels underscore some of the most striking differences between diffusion and wave phenomena. Thus, we review the kernels of both the diffusion and wave equations to clarify how they differ. For simplicity, we focus on the one-dimensional case in the reminder of this section. We also provide fundamental theorems from an energetic perspective that characterize each dynamic in Appendix A.
@@
We recall the fundamental solution (Green’s function) of the one-dimensional diffusion equation evans2022partial with the initial condition $x(u,0)=x_{0}(u)$. Its kernel (Green’s function) on $\mathbb{R}$ is given by
$G(u,t)=\frac{1}{\sqrt{4\pi t}}\exp\!\biggl{(}-\frac{u^{2}}{4t}\biggr{)},\quad t>0,$ (5)
where we ignore the diffusive coefficient $D$. Accordingly, the solution with initial data $x_{0}$ can be written in convolution form: $x(u,t)=\int_{-\infty}^{\infty}G(u-v,\,t)\,x_{0}(v)\,\mathrm{d}v.$ Key properties of $G$ in one dimension are as follows: Firstly, $G(u,t)>0$ for all $u\in\mathbb{R}$ and decays exponentially as $|u|\to\infty$. And secondly, the heat kernel smooths out irregularities in the initial data, which is referred to as the smoothing effect. The first property shows that, while the diffusive influence extends over all space, it becomes exponentially small far from the origin. The second property implies that diffusion rapidly weakens high-frequency components, inherently suppressing local structural details.
@@
Next, we consider the one-dimensional wave equation with initial conditions $x(u,0)=f(u)$ and $\frac{\partial x}{\partial t}(u,0)=g(u)$. By d’Alembert’s formula pinchover2005introduction , the general solution is
$x(u,t)=\frac{1}{2}\bigl{[}f(u-v\,t)+f(u+v\,t)\bigr{]}+\frac{1}{2v}\int_{u-v\,t}^{\,u+v\,t}g(s)\,\mathrm{d}s,$ (6)
which represents left- and right-traveling waves emanating from the initial data. Key properties of this solution are as follows: Firstly, under wave propagation, strong singularities in the initial data can persist. Thus, the signal profile propagates at finite speed, rather than being smoothed out everywhere. And secondly, the general solution reflects conservation of energy: waves transport energy through the domain without diffusive spreading. Hence, in contrast to the diffusion equation, the wave equation exhibits a sharp, finite wavefront traveling at speed $v$. This difference in how features propagate is closely related to the fact that the wave equation conserves energy, whereas the diffusion equation dissipates it. The energy behaviors of each dynamics are shown in Appendix A.
@@ 3 Wavy Transformer
Figure 1: Schematic of Wavy Transformer block, combining wavy attention layers and velocity- specific layer-normalization and feed-forward layers. Each layer is designed to preserve state-velocity relationship. Post-LN is assumed here; the Pre-LN case is discussed in Appendix B.2 xiong2020layer .
The above discussion and the fundamental theorems 1 and 2 in Appendix A indicate that the dynamics governed by the diffusion equation (3) converge to a constant uniform state, whereas those governed by the wave equation (4) oscillates while conserving its energy without net dissipation. However, conventional transformers implicitly rely on the dissipative diffusion dynamics, causing their hidden states to converge toward a uniformly constant state. Therefore, before presenting the architecture of Wavy Transformer, we first show that the implicit dynamics of attention layers can be interpreted as graph neural diffusion chamberlain2021grand on a complete graph.
@@ 3.1 Attention as Graph Neural Diffusion
for $l=1,\cdots,N$, where $N$ is the number of layers and $\tau>0$ is a fixed time interval parameter. Also, $\mathbf{X}^{l}$ denotes the hidden states at time $l\tau$. Interestingly, if we consider $\tau=\frac{1}{2}$, we can get $\mathbf{X}^{l+1}=\frac{1}{2}(\mathbf{A}\mathbf{X}^{l}+\mathbf{X}^{l})$. Usually, since each attention layer is followed by a layer normalization (LN), the scale transformation can be ignored. Therefore, this discrete diffusion equation is essentially identical to the attention layer update $\mathbf{X}^{l+1}=\mathbf{A}\mathbf{X}^{l}+\mathbf{X}^{l}$. Although we here assume Post-LN xiong2020layer , in the same way, the attention with Pre-LN can be understood as diffusion–reaction equation, which is explained in Appendix B. In this sense, the dynamics of attention layers can be understood as a graph diffusion on a complete graph. Furthermore, because diffusion decreases the system’s energy (as shown in Appendix A), this interpretation aligns with the discussion in wu2023difformer ; nguyen2023mitigating .
As a result, the conventional transformer architecture can be viewed as executing a diffusion process on a complete graph in which each node represents a token embedding. Moreover, the attention matrix $\mathbf{A}$ defines anisotropic diffusivity interactions among tokens. This perspective suggests that the over-smoothing observed in transformers results from the dissipative smoothing effect of a diffusion equation, analogous to heat flow. An intuitive explanation of this phenomenon is provided in Appendix C. By viewing diffusion as a process that smooths small-scale fluctuation erasing high-frequency components, it becomes clear why the attention layer acts as a low-pass filter wang2022antioversmooth . It also agrees with the interpretation of GNNs as performing low-pass filtering, as noted in nt2019revisitinggraphneuralnetworks .
@@ 3.2 Wavy Dynamics Based Attention
We introduce a novel transformer block based on wavy rather than diffusive dynamics, motivated by the implicit reliance of conventional transformers on diffusive dynamics and some discussion for GNNs eliasof2021pde ; graphcon . Figure 1 shows a schematic of Wavy Transformer block, which includes a wavy attention layer, velocity-specific layer normalization, and a velocity-oriented feed-forward network. We begin by considering the wave equation on a complete graph, analogous to the graph neural diffusion chamberlain2021grand :
@@ 4 Related Works
4.1 Mitigating Over-Smoothing: Existing Approaches and This Study
@@ 4.2 Baselines for Comparative Study
In our comparative study, we include BERT and DeiT as representative baselines for NLP and CV applications, respectively.
BERT: Bidirectional Encoder Representations from Transformers
BERT, introduced by Devlin et al. devlin-etal-2019-bert , is a deep bidirectional transformer encoder trained with masked language modeling and next sentence prediction objectives. By conditioning on both left and right context in every layer, BERT learns rich contextual embeddings that serve as a standard benchmark for a wide range of NLP tasks. Thus, we chose BERT as the canonical NLP model for comparison.
DeiT: Data-efficient Image Transformers
DeiT, proposed by Touvron et al. DeiT , adapts the transformer architecture to vision tasks without massive pre-training. DeiT exemplifies a state-of-the-art CV transformer, which we use as our baseline model for CV task. Also, we review an existing over-smoothing mitigation technique wang2022antioversmooth : FeatScale, which re-weights features to boost high-frequency signals. We then show Wavy Transformer can seamlessly integrate with it to enhance their performance.
@@ 5 Experiments
In this section, we present a comprehensive experimental evaluation of Wavy Transformer block for NLP tasks under a BERT-like pretraining framework devlin-etal-2019-bert , measuring (i) pretraining perplexity and MLM accuracy, (ii) fine-tuning performance on GLUE tasks wang2018glue , and (iii) over-smoothing behavior. Moreover, to demonstrate the generality and effectiveness of our approach with other types of transformer-based backbones DeiT DeiT , we validate Wavy Transformer on representative computer vision tasks, specifically ImageNet classification imagenet_cvpr09 . Experiments used servers with eight V100 GPUs, four L40 GPUs, or four H100 GPUs, selected according to task size.

### v2

@@ Abstract
Transformers have achieved remarkable success across natural language processing (NLP) and computer vision (CV). However, deep transformer models often suffer from an over-smoothing issue, in which token representations converge to similar values as they pass through successive transformer blocks. In this paper, we establish an equivalence between the hidden-state dynamics induced by stacked attention layers and graph neural diffusion on a complete graph. From this perspective, over-smoothing can be interpreted as a consequence of the dissipative nature of the underlying diffusion dynamics. Motivated by this physical interpretation, we propose Wavy Transformer, which consists of a novel attention layer based on second-order wavy dynamics. We also introduce a feed-forward network and a normalization layer designed to preserve the physical state-velocity relationship under the chain rule, thereby extending the transformer architecture. We further validate our proposed techniques on various transformer models for NLP, CV, and sparse-graph tasks. The results consistently demonstrate that Wavy Transformer improves performance with minimal additional parameters and no extra hyperparameter tuning. Source code and models are available at https://github.com/noguchisatoshi/Wavy-Transformer.
@@ 1 Introduction
Based on the above discussion, in this paper, we examine over-smoothing in transformers from the perspective of physical dynamical systems. In addition, we propose a new type of attention layer inspired by physical dynamical systems. Our approach originates from interpreting the dynamics of hidden states induced by stacked attention layers as graph neural diffusion chamberlain2021grand on a complete graph. From this viewpoint, over-smoothing can be seen as a consequence of the dissipative nature of the underlying diffusive dynamics. Motivated by this physical interpretation, we introduce Wavy Transformer, which consists of a novel attention layer based on second-order wavy dynamics. The energy-preserving property and the oscillatory behavior of the wave equation are expected to help mitigate the over-smoothing problems in transformers. We also introduce a physically inspired feed-forward network and a normalization layer designed to preserve the state-velocity relationship under the chain rule. By combining these components, we build Wavy Transformer to extend the conventional transformer architecture. Wavy Transformer block is easy to use and can be integrated into various transformer-like architecture. Furthermore, to demonstrate the effectiveness and generality of Wavy Transformer, we conduct extensive experiments on NLP, CV, and sparse-graph tasks.
@@ 2.1 Attention and Over-smoothing
Despite their remarkable performance across a wide range of NLP and CV applications, deep transformer models often suffer from an over‑smoothing problem: as the network deepens, token representations converge and become indistinguishable from one another shi2022revisiting ; wang2022antioversmooth ; nguyen2023mitigating . As a result, deeper transformers do not always outperform their shallower counterparts zhou2021deepvit . From a theoretical viewpoint, the attention layer can be understood as a low‑pass filter wang2022antioversmooth , continuously erasing high‑frequency information and thus reducing feature expressiveness in deeper layers. Consequently, this tendency toward over‑smoothing of stacked the attention layers in deep transformer presents a critical challenge that motivates our study.
@@ 2.2 Fundamental Dynamics on Manifolds
$\frac{\partial x(s,t)}{\partial t}=\mathrm{div}\bigl(D\,\nabla x(s,t)\bigr)=D\,\Delta x(s,t),$ (3)
where $s\in M$ denotes a point (spatial coordinate) on the manifold, $t\in[0,T]$ denotes time, $D\geq 0$ is a diffusivity coefficient, $\nabla$ and $\mathrm{div}$ are respectively the gradient and divergence operators induced by the Riemannian metric on $M$, and $\Delta=\mathrm{div}\circ\nabla$ is the Laplace–Beltrami operator. Here, $D$ is assumed to be constant for simplicity. In physics, Eq.(3) models processes such as heat conduction or particle diffusion. In contrast, the hyperbolic (wave) equation on $M$ takes the form
$\frac{\partial^{2}x(s,t)}{\partial t^{2}}=\mathrm{div}\bigl(v^{2}\,\nabla x(s,t)\bigr)=v^{2}\,\Delta x(s,t),$ (4)
where $v\geq 0$ is a constant wave‑propagation speed. Equation (4) governs wave-like phenomena in continuous media. We highlight the key differences between the dynamics of these PDEs.
@@ 2.2.1 Kernel Perspective of Diffusion and Wave Equations
Fundamental solutions, which are often called "kernels", represent the response to an initial impulse (e.g., a Dirac delta function). These kernels underscore some of the most striking differences between diffusion and wave phenomena. Thus, we review the kernels of both the diffusion and wave equations to clarify how they differ. For simplicity, we focus on the one-dimensional case in the reminder of this section. We also provide fundamental theorems from an energetic perspective that characterize each dynamic in Appendix C.
@@
We recall the fundamental solution (Green’s function) of the one‑dimensional diffusion equation evans2022partial with the initial condition $x(s,0)=x_{0}(s)$. Its kernel (Green’s function) on $\mathbb{R}$ is given by
$G(s,t)=\frac{1}{\sqrt{4\pi t}}\exp\!\biggl(-\frac{s^{2}}{4t}\biggr),\quad t>0,$ (5)
where we ignore the diffusive coefficient $D$. Accordingly, the solution with initial data $x_{0}$ can be written in convolution form: $x(s,t)=\int_{-\infty}^{\infty}G(s-v,\,t)\,x_{0}(v)\,\mathrm{d}v.$ Key properties of $G$ in one dimension are as follows: Firstly, $G(s,t)>0$ for all $s\in\mathbb{R}$ and decays exponentially as $|s|\to\infty$. And secondly, the heat kernel smooths out irregularities in the initial data, which is referred to as the smoothing effect. The first property shows that, while the diffusive influence extends over all space, it becomes exponentially small far from the origin. The second property implies that diffusion rapidly weakens high‑frequency components, inherently suppressing local structural details.
@@
Next, we consider the one‑dimensional wave equation with initial conditions $x(s,0)=f(s)$ and $\frac{\partial x}{\partial t}(s,0)=g(s)$. By d’Alembert’s formula pinchover2005introduction , the general solution is
$x(s,t)=\frac{1}{2}\bigl[f(s-v\,t)+f(s+v\,t)\bigr]+\frac{1}{2v}\int_{s-v\,t}^{\,s+v\,t}g(r)\,\mathrm{d}r,$ (6)
which represents left‑ and right‑traveling waves emanating from the initial data. Key properties of this solution are as follows: Firstly, under wave propagation, strong singularities in the initial data can persist. Thus, the signal profile propagates at finite speed, rather than being smoothed out everywhere. And secondly, the general solution reflects conservation of energy: waves transport energy through the domain without diffusive spreading. Hence, in contrast to the diffusion equation, the wave equation exhibits a sharp, finite wavefront traveling at speed $v$. This difference in how features propagate is closely related to the fact that the wave equation conserves energy, whereas the diffusion equation dissipates it. The energy behaviors of each dynamics are shown in Appendix C.
@@ 3 Wavy Transformer
Figure 1: Schematic of Wavy Transformer block, combining wavy attention layers and velocity‑ specific layer‑normalization and feed‑forward layers. Each layer is designed to preserve state-velocity relationship. Post-LN is assumed here; the Pre-LN case is discussed in Appendix D.2 xiong2020layer .
The above discussion and the fundamental theorems 1 and 2 in Appendix C indicate that the dynamics governed by the diffusion equation (3) converge to a constant uniform state, whereas those governed by the wave equation (4) oscillates while conserving its energy without net dissipation. However, conventional transformers implicitly rely on the dissipative diffusion dynamics, causing their hidden states to converge toward a uniformly constant state. Therefore, before presenting the architecture of Wavy Transformer, we first show that the implicit dynamics of attention layers can be interpreted as graph neural diffusion chamberlain2021grand on a complete graph.
@@ 3.1 Attention as Graph Neural Diffusion
for $l=1,\cdots,N$, where $N$ is the number of layers and $\tau>0$ is a fixed time interval parameter. Also, $\mathbf{X}^{l}$ denotes the hidden states at time $l\tau$. Interestingly, if we consider $\tau=\frac{1}{2}$, we can get $\mathbf{X}^{l+1}=\frac{1}{2}(\mathbf{A}\mathbf{X}^{l}+\mathbf{X}^{l})$. Usually, since each attention layer is followed by a layer normalization (LN), the scale transformation can be ignored. Therefore, this discrete diffusion equation is essentially identical to the attention layer update $\mathbf{X}^{l+1}=\mathbf{A}\mathbf{X}^{l}+\mathbf{X}^{l}$. Although we here assume Post-LN xiong2020layer , in the same way, the attention with Pre-LN can be understood as diffusion–reaction equation, which is explained in Appendix D. In this sense, the dynamics of attention layers can be understood as a graph diffusion on a complete graph. Furthermore, because diffusion decreases the system’s energy (as shown in Appendix C), this interpretation aligns with the discussion in wu2023difformer ; nguyen2023mitigating .
As a result, the conventional transformer architecture can be viewed as executing a diffusion process on a complete graph in which each node represents a token embedding. Moreover, the attention matrix $\mathbf{A}$ defines anisotropic diffusivity interactions among tokens. This perspective suggests that the over-smoothing observed in transformers results from the dissipative smoothing effect of a diffusion equation, analogous to heat flow. An intuitive explanation of this phenomenon is provided in Appendix E. By viewing diffusion as a process that smooths small-scale fluctuation erasing high‑frequency components, it becomes clear why the attention layer acts as a low‑pass filter wang2022antioversmooth . It also agrees with the interpretation of GNNs as performing low‑pass filtering, as noted in nt2019revisitinggraphneuralnetworks .
@@ 3.2 Wavy Dynamics Based Attention
We introduce a novel transformer block based on wavy rather than diffusive dynamics, motivated by the implicit reliance of conventional transformers on diffusive dynamics and some discussion for GNNs eliasof2021pde ; graphcon . Figure 1 shows a schematic of Wavy Transformer block, which includes a wavy attention layer, velocity‑specific layer normalization, and a velocity‑oriented feed‑forward network. We begin by considering the wave equation on a complete graph, analogous to the graph neural diffusion chamberlain2021grand :
@@ 4 Related Works
4.1 Mitigating Over‑Smoothing: Existing Approaches and This Study
@@
4.2 Physics-Inspired GNN and Transformers
Our study is complementary to prior physics-inspired GNNs and transformers. Graph-CON graphcon and PDE-GCN eliasof2021pde introduce oscillatory or PDE-motivated updates on sparse graphs; in contrast, we target complete-graph attention in transformers. Relatedly, Gravina et al.gravina2025swan analyze over-squashing in sparse-graph GNNs, whereas we focus on over-smoothing in all-to-all attention. Deng et al.deng2025denoising introduce Hamiltonian structure into transformers through a loss, while we directly replace the residual dynamics with a second-order wave update without auxiliary losses (see also HNNs NEURIPS2019_26cd8eca ).
@@ 4.3 Baselines for Comparative Study
In our comparative study, we include BERT, DeiT, and DIFFormer as representative baselines for NLP, CV, and sparse-graph applications, respectively.
BERT: Bidirectional Encoder Representations from Transformers BERT, introduced by Devlin et al. devlin-etal-2019-bert , is a deep bidirectional transformer encoder trained with masked language modeling and next sentence prediction objectives. By conditioning on both left and right context in every layer, BERT learns rich contextual embeddings that serve as a standard benchmark for a wide range of NLP tasks. Thus, we chose BERT as the canonical NLP model for comparison.
DeiT: Data‑efficient Image Transformers DeiT, proposed by Touvron et al. DeiT , adapts the transformer architecture to vision tasks without massive pre‑training. DeiT exemplifies a state‑of‑the‑art CV transformer, which we use as our baseline model for CV task. Also, we review an existing over‑smoothing mitigation technique wang2022antioversmooth : FeatScale, which re‑weights features to boost high‑frequency signals. We then show Wavy Transformer can seamlessly integrate with it to enhance their performance.
DIFFormer: Diffusion-based Graph Transformers As a representative graph transformer, we adopt DIFFormer wu2023difformer as our diffusion baseline on sparse-graph benchmarks.
@@ 5 Experiments
In this section, we present a comprehensive experimental evaluation of Wavy Transformer block for NLP tasks under a BERT-like pretraining framework devlin-etal-2019-bert , measuring (i) pretraining perplexity (PPL) and MLM accuracy, (ii) fine-tuning performance on GLUE tasks wang2018glue , and (iii) over-smoothing behavior. Moreover, to demonstrate the generality and effectiveness of our approach with other types of transformer-based backbones DeiT DeiT , we validate Wavy Transformer on representative computer vision tasks, specifically ImageNet classification imagenet_cvpr09 . Also, we evaluated the performance of Wavy Transformer on sparse-graph benchmark based on DIFFormer wu2023difformer . Experiments used servers with eight V100 GPUs, four L40 GPUs, or four H100 GPUs, selected according to task size.
Model Variants
We evaluate two variants of the Wavy Transformer. Both share the same attention backbone; they differ only in how the velocity residual is realized. Full Wave. A second-order residual realized with an explicit velocity branch that includes a FFN and a LN on the velocity path as shown in Figures 1, yielding stronger physical constraints at additional computational costs. Light Wave. A momentum-only realization that keeps the second-order effect while removing FFN/LN for velocity. The update is $\mathbf{X}^{l+1}=\tau\mathbf{A}\mathbf{X}^{l}+(1-\tau)\mathbf{X}^{l}+\boldsymbol{\lambda}\big(\mathbf{X}^{l}-\mathbf{X}^{l-1}\big),$ where $\boldsymbol{\lambda}\!\in\![0,1]^{d}$ controls the mix between the diffusion part and the wave (momentum) part. Ignoring the coefficient mismatch between the diffusion term ($\tau$) and the momentum term ($\tau^{2}$), setting $\boldsymbol{\lambda}=\mathbf{0}$ recovers diffusion-only, whereas $\boldsymbol{\lambda}=\mathbf{1}$ corresponds to the wave-only residual. Formal derivations appear in Appendix. B.